OpenStudy (anonymous):

is that set in the subspaces of R3 {x,y,z | x+y+z=-7

OpenStudy (turingtest):

\[\vec u=x+y+z=-7\]\[\vec v= a+b+c=-7\]\[\vec u+\vec v=?\]

OpenStudy (turingtest):

\[\vec u=x+y+z=-7\]\[c\vec u=cx+cy+cz=-c7;c\neq0,1\]

OpenStudy (turingtest):

are either of these in the subspace?

OpenStudy (anonymous):

no

OpenStudy (turingtest):

then you have your answer :) Welcome to Open Study!

OpenStudy (anonymous):

i c, thanks alot